# Spin-resolved nonlocal transport in proximitized Rashba nanowires

### v1

@@ I Introduction
Recent state-of-the-art experiments have demonstrated unprecedented control over nanofabrication processes Shabani_2016 ; Zhang2017 ; Gazibegovic_2017 ; Thomas_2019 ; Mayer2020 ; Heedt_2021 ; Moehle2021 ; OConnell_2021 ; Badawy_2024 and precise tuning of the parameters of such devices Mazur_2022 ; Loo_2023 ; Tunable_car_PRX . One of the most accessible experimental methods to investigate and study properties of quasiparticles hosted in such hybrid structures are quantum transport techniques. The local transport spectroscopy techniques can provide some insight about localized in-gap quasiparticle states e.g zero energy Majorana fermions or non-zero energy ABSs via local conductance peaks as predicted by theory Law2009 ; Akhmerov_2009 ; Sau_2010 ; Flensberg2010 ; Wimmer_2011 ; Fidkowski_2012 ; Diego_2013 ; Chevallier_2016 and tested experimentally Sasaki_2011 ; Mourik2012 ; Das2012 ; Deng_2012 ; Churchill_2013 ; Deng_2016 ; Nichele_2017 ; Lutchyn_2018 ; Anselmetti_2019 ; Vaitiekenas_2020 ; Schneider_2021 ; Wang_2022 . However, one has to keep in mind that such local transport measurements can be inconclusive in detecting topological superconducting phases since the observed zero-bias peaks can have origins that are different from Majorana fermions Kells_2012 ; Lee_2012 ; Lee_Nat_Nan_2014 ; Cayao_2015 ; Ptok_2017 ; Liu_2017 ; Reeg_2018 ; Penaranda2018 ; Moore2018 ; Vuik2019 ; Woods2019 ; Liu2019 ; Chen2019 ; Pan_2020 ; Alspaugh2020 ; Junger2020 ; Valentini2021 ; Yu_2021 ; Hess_2021_sc ; Hess_2023 . On the other hand, multiterminal transport techniques can provide additional information about the system such as superconducting energy gap anisotropy, nonlocal nature of the quasiparticles, their effective charge associated with electron and hole composition, induced gap closing, and, potentially, about the TPT Flatte_1995 ; Wohlman_2008 ; Schindele_2014 ; Schindele_2014 ; Lobos_2015 ; Gramich_2017 ; Rosdahl_2018 ; Zhang_2019 ; Nonlocal_BCS_charge_th ; Melo_2021 ; Pikulin_arxiv ; Pan_2021 ; Hess_2021 ; Menard_2020 ; Puglia_2021 ; Poschl_2022 ; Dourado_arxiv ; Hess_2023 ; topological_gap ; Thamm_2024 .
@@
Moreover, we identify parameter regimes for which the quasiparticle charge is nearly zero, indicating an equal amount of electron and hole components even though the system undergoes band inversion. In such a regime, the sign of nonlocal conductance related to the quasiparticle spin are particularly robust to onsite disorder. For completeness, we also calculate the local conductance, which also depends on the spin polarization of the probed quasiparticle states, obtaining results consistent with the spin-selective Andreev reflection process SSAR ; SSAR2_ex ; SSAR2_th ; SSAR_Dom ; Jeon_science_2017 ; Jeon_prb_2018 ; Wang_prl_2021 . On the other hand, in setups with two normal leads, we obtain a very good mapping between the sign of nonlocal conductance and the sign of quasiparticle charge as expected theoretically Nonlocal_BCS_charge_th ; Tunable_car_PRL_theory and verified experimentally Tunable_car_PRX .
Our findings describe a general phenomenon and can be used to precisely detect the local spin density of quasiparticles not only in proximitized Rashba nanowires but also in other systems hosting spin-polarized quasiparticle states such as Yu-Shiba-Rusinov states Yu_1965 ; Shiba_1968 ; Rusinov_1969 ; Yazdani_1997 , spin chains Choy_2011 ; Martin_2012 ; Yazdani_2013 ; Pientka ; Klinovaja_2013 ; Braunecker_2013 ; Vazifeh_2013 ; Yazdani_2014 ; Ruby_2015 ; Pawlak_2016 ; Schneider_2022 ; Kuster_2022 , Andreev spin qubits Andreev_sq_th ; Park_2017 ; Mishra_2021 ; Andreev_sq_exp ; Vidal_2023 ; Spectr_Spin_Andreec_SC ; Hoffman_2024 , Caroli-de Gennes-Matricorn and Majorana vortex states Volovik_1999 ; Caroli1964 ; Xu2015 ; Sun2016 ; Chiu2020 etc. Furthermore, our results offer insights for designing devices functioning as Cooper pair splitters and for realizing poor man’s Majorana fermions in minimal Kitaev chain models in which precise tunability over ECT and CAR processes is crucial Tunable_car_PRX ; MKC_ex ; MKC_ex_3dots . In our simulations, we employ realistic parameters. It is noteworthy that the coupling of side-leads to nanowires has already been experimentally demonstrated, both for normal leads topological_gap , for quantum dots Lee_Nat_Nan_2014 ; spin_filter_delftNC ; spin_filter_dots_Marcus , and, notably, for ferromagnetic leads Andreas_spin_ent_N . Hence, our predictions are directly amenable to experimental verification.
@@ II Model
In order to investigate the transport properties of the system, we employ the Blonder-Tinkham-Klapwijk formalism BTK in which the zero-temperature nonlocal conductance $G_{ij}(E)$ between leads $i$ and $j$ is given by
@@
The corresponding band structure of the leads $E_{L,R}(k)$ as a function of momentum $k$ is schematically depicted on Fig. 2 together with the band structure of the proximitized Rashba nanowire $E_{RNW}(k)$ and quasiparticle spin $S_{x}(k)=\Phi^{\dagger}(k)\tau_{z}\sigma_{x}\Phi(k)$ (see App. A for details). Again, the spin quantization axis is assumed to be in the $x$ direction, along the applied magnetic field. The leads are tunnel-coupled to the nanowire at the positions $x_{L}$ and $x_{R}$, respectively, for the left and the right lead with positive hopping amplitude $t_{\Gamma}(<t)$.
Figure 2: Energy bands of the left spin up polarized lead $E_{L}(k)$, the proximitized Rashba nanowire $E_{RNW}(k)$ and the right lead which is either (a) spin-down or (b) spin-up polarized or (c) normal $E_{R}(k)$. The $x$ component (along B-field) of the spin polarization of quasiparticles from the Andreev band, $S_{x}(k)$, is indicated by the colorbar. Shaded gray areas mark an energy window $\Delta_{i}<E<\Delta_{ex}$ in which spin sensitive transport occurs. Among many possible scattering processes, we denote the dominant ones contributing to the nonlocal conductances (b) $G_{L(+)R(-)}$, (c) $G_{L(+)R(+)}$, and (d) $G_{L(+)R(N)}$, which are either CAR (negative conductance) or ECT (positive conductance) processes. The strength of the corresponding signal (high or low) is reflected in the line thickness. Empty circles denote holes while filled ones denote electrons and half-filled ones correspond to quasiparticles in the nanowire. We set the following parameters for the nanowire: $\mu=0$, $t=0.27$ meV, $\Delta_{sc}=0.25$ meV, $\Delta_{Z}=0.2$ meV (nontopological phase), $\alpha=50$ meVnm ($E_{\rm SO}=0.23$ meV). For the spin up (down) polarized leads, we set $\mu_{\rm lead}=0$ and $M_{x}=-0.3$ meV ($M_{x}=0.3$ meV) and, for the normal lead, we set $\mu_{\rm lead}=0.3$ meV and $M_{x}=0$.
@@ III.1 Schematic picture of nonlocal transport
We present first a schematic (physical) picture of the dominant processes contributing to the nonlocal conductance $G_{LR}$ for three different configurations of the leads, which is then followed by the presentation of exact numerical results in subsequent sections. Regarding the proximitized Rashba nanowire, the extended Andreev band quasiparticles with energies $E_{RNW}(k)<\Delta_{sc}$ have a certain sign of the spin if the exterior gap exceeds the interior one, $\Delta_{i}<\Delta_{ex}$, which for $\mu=0$ is satisfied if $0<\Delta_{Z}<2\Delta_{sc}$. In the topologically trivial state (here, $\Delta_{Z}<\Delta_{sc}$ for $\mu=0$), the quasiparticle states with the lowest positive energy around $k\approx 0$ have negative spin polarization $S_{x}(k)$ as illustrated on Fig. 2(a)-(c) (middle panels).
When the left lead is spin-up (+) and the right lead spin-down polarized (-), the spin of the injected charge carrier with positive energy matches the negative spin of quasiparticles in the proximitized region. However, due to spin mismatch it cannot enter the left lead as an electron and has to be transformed into hole which is spin-down polarized (see Fig. 2(a) for details). This leads to a strong CAR dominated nonlocal conductance $G_{L(+)R(-)}$ signal characterized by a negative sign.
@@
Finally, we consider the case with the left lead being spin-up polarized while the right lead is normal (unpolarized). Due to spin degeneracy in the right lead, the electron can freely enter the proximitized region regardless of the quasiparticle spin polarization and enter the left lead as spin-down (up) hole (electron). This supports strong CAR (ECT) dominated nonlocal conductances for positive (negative) energies (see Fig. 2(c) for details). In the following sections, we support the picture presented here by exact numerical simulations of the spin-dependent quantum transport.
@@ III.2 Quasiparticle spin detection
We start the discussion with presenting results from transport simulations for systems with two leads that have the same spin polarization either up (+) or down (-). We observe a strong positive nonlocal conductance $G_{L(\pm)R(\pm)}(\Delta_{Z},E_{F})=G_{R(\pm)L(\pm)}(\Delta_{Z},E_{F})$, if the spin of the probed ABS is the same as the spin of the electrons in the leads. In this case the transport is dominated by ECT processes. For example, when the leads are spin up (spin down) polarized one can observe strong positive nonlocal conductance signals as depicted on Fig. 3(b’) [Fig. 3(b”)] for energies corresponding to the lowest bulk states and values of the Zeeman energies for which corresponding quasiparticle states also have positive (negative) sign of the local spin density at the positions of the leads. The
@@
appearance of weak negative nonlocal conductance signals (CAR) is due to the presence of SOI in the nanowire that allows for normally forbidden spin-flip transport processes. On the other hand, by analyzing plots of the local conductance shown in Fig.  3(c’) and (c”), we notice that conductance peaks are more (less) pronounced when the spin polarization in the leads matches (does not match) the spin polarization of the probed quasiparticles. This observation is consistent with so-called SSAR processes SSAR ; SSAR2_ex ; SSAR2_th ; SSAR_Dom ; Jeon_science_2017 ; Jeon_prb_2018 ; Wang_prl_2021 . We obtain analogous results with ECT dominated signals when the chemical potential deviates from $\mu=0$ while $\Delta_{Z}$ is fixed (see Fig. 4). In this case, there are two TPTs at $\mu_{c}^{\pm}=\pm\sqrt{\Delta_{Z}^{2}-\Delta_{sc}^{2}}$. Still, we again can identify these TPTs by simply looking at the sign of the nonlocal conductances.
@@
Figure 6: The same as for Fig. 5 but now the chemical potential $\mu$ is varied and the Zeeman energy is kept fixed to $\Delta_{Z}=0.4$ meV. Here, we can see again that the transport is dominated by CAR processes (negative conductance), which can be controled by tuning $\mu$ (e.g. via backgate).
@@
We have also considered the case when the left lead is spin polarized and the right one is normal - spin unpolarized (spin degenerate). In such a setup, the nonlocal conductance $G_{L(\pm)R(N)}$ has a strong positive (negative) value when the polarization of the left lead is the same as (opposite to) the spin polarization of the low-energy quasiparticles from the Andreev band. Such gate configurations make the system an ideal platform for detecting the sign of the local spin polarization of quasiparticles, which is directly linked to the sign of the nonlocal conductance. Therefore, such setups are very suitable for the detection of spin inversion of the one-dimensional Andreev bands induced by the TPT. Furthermore, such setups are optimal for switching between ECT and CAR processes in Rashba nanowires either by tuning the magnetic field [see Fig. 7(a,b)] or the chemical potential [see Fig. 8(a,b)]. For this case, we have also calculated $G_{L(\pm)R(N)}(x_{L})$ as a function of the position of the left lead while the position of the right one is fixed. We plot it together with quasiparticle spin and charge densities for the system in the topological as well as nontopological phase, see Fig. 9. We note again that the sign of the nonlocal conductance is correlated with the sign of the local spin density. We also note that when the leads are attached to the ends of the system, the sign of the nonlocal conductance does not change when the system undergoes a TPT since the local spin density has the same sign at the ends of the nanowire (see Fig. 9 for details) in both topological and nontopological phases. Here, on purpose, we choose shorter nanowires $L=3\ \mu$m to illustrate clearly the case when the sign of the quasiparticle spin density $S_{x}(x)$ is position dependent [see orange curve on Fig. 9(b)]. As a consequence, in this case the sign of the nonlocal conductances $G_{L(\pm)R(N)}(x_{L})$ depends on the position of the left lead [see solid and dashed black curves on Fig. 9(b)].
@@
Finally, we show that our results are robust against disorder, by adding random on-site fluctuations to the onsite chemical potential $\mu_{i}=\mu+\delta\mu_{i}$ in Eq. (1) with $|\delta\mu_{i}|\leq\Delta_{sc}$ as shown on Fig. 10(d). One can see that the presence of disorder affects the energy levels [see Fig. 10(a)] compared to the clean system [see Fig. 3(a)], however, the sign of the local spin density is very robust to such disorder [see Fig 10 (a,e)]. Moreover, as a consequence, our transport study shows that the sign of the nonlocal conductance is also not affected by the disorder [see Fig. 5(a, b’), Fig. 6(a,b’), Fig. 5(a,b’), Fig. 10(b,b’,b”,c,c’,c”)]. Importantly, in the charge neutrality regime where $\mu=0$ and where the band inversion is driven by the Zeeman energy, the spin-dependent nonlocal conductances are clearly more robust to the onsite disorder [see Fig. 5(a’,b’), Fig. 10(b,b’, b”,c,c’,c”)] than in the case when $\mu\neq 0$ [see Fig. 6(a’,b’)].

### v2

@@ I Introduction
Recent state-of-the-art experiments have demonstrated unprecedented control over nanofabrication processes Shabani_2016 ; Zhang2017 ; Gazibegovic_2017 ; Thomas_2019 ; Mayer2020 ; Heedt_2021 ; Moehle2021 ; OConnell_2021 ; Badawy_2024 and precise tuning of the parameters of such devices Mazur_2022 ; Loo_2023 ; Tunable_car_PRX . One of the most accessible experimental methods to investigate and study properties of quasiparticles hosted in such hybrid structures are quantum transport techniques. The local transport spectroscopy techniques can provide some insight about localized in-gap quasiparticle states e.g zero energy Majorana fermions or non-zero energy ABSs via local conductance peaks as predicted by theory Law2009 ; Akhmerov_2009 ; Sau_2010 ; Flensberg2010 ; Wimmer_2011 ; Fidkowski_2012 ; Diego_2013 ; Chevallier_2016 and tested experimentally Sasaki_2011 ; Mourik2012 ; Das2012 ; Deng_2012 ; Churchill_2013 ; Deng_2016 ; Nichele_2017 ; Lutchyn_2018 ; Anselmetti_2019 ; Vaitiekenas_2020 ; Schneider_2022 ; Wang_2022 . However, one has to keep in mind that such local transport measurements can be inconclusive in detecting topological superconducting phases since the observed zero-bias peaks can have origins that are different from Majorana fermions Kells_2012 ; Lee_2012 ; Lee_Nat_Nan_2014 ; Cayao_2015 ; Ptok_2017 ; Liu_2017 ; Reeg_2018 ; Penaranda2018 ; Moore2018 ; Vuik2019 ; Woods2019 ; Liu2019 ; Chen2019 ; Pan_2020 ; Alspaugh2020 ; Junger2020 ; Valentini2021 ; Yu_2021 ; Hess_2021_sc ; Hess_2023 . On the other hand, multiterminal transport techniques can provide additional information about the system such as superconducting energy gap anisotropy, nonlocal nature of the quasiparticles, their effective charge associated with electron and hole composition, induced gap closing, and, potentially, about the TPT Flatte_1995 ; Wohlman_2008 ; Schindele_2014 ; Schindele_2014 ; Lobos_2015 ; Gramich_2017 ; Rosdahl_2018 ; Zhang_2019 ; Nonlocal_BCS_charge_th ; Melo_2021 ; Pikulin_arxiv ; Pan_2021 ; Hess_2021 ; Menard_2020 ; Puglia_2021 ; Poschl_2022 ; Dourado_arxiv ; Hess_2023 ; topological_gap ; Thamm_2024 .
@@
Moreover, we identify parameter regimes for which the quasiparticle charge is nearly zero, indicating an equal amount of electron and hole components even though the system undergoes band inversion. In such a regime, the sign of nonlocal conductance related to the quasiparticle spin are particularly robust to onsite disorder. For completeness, we also calculate the local conductance, which also depends on the spin polarization of the probed quasiparticle states, obtaining results consistent with the spin-selective Andreev reflection process SSAR ; SSAR_ ; SSAR2_ex ; SSAR2_th ; SSAR_Dom ; Jeon_science_2017 ; Jeon_prb_2018 ; Wang_prl_2021 ; Karan_2024 . On the other hand, in setups with two normal leads, we obtain a very good mapping between the sign of nonlocal conductance and the sign of quasiparticle charge as expected theoretically Nonlocal_BCS_charge_th ; Tunable_car_PRL_theory and verified experimentally Tunable_car_PRX .
Our findings describe a general phenomenon and can be used to precisely detect the local spin density of quasiparticles not only in proximitized Rashba nanowires but also in other systems hosting spin-polarized quasiparticle states such as Yu-Shiba-Rusinov states (YSR) Yu_1965 ; Shiba_1968 ; Rusinov_1969 ; Yazdani_1997 , spin chains Schneider_2022 ; Choy_2011 ; Martin_2012 ; Yazdani_2013 ; Pientka ; Klinovaja_2013 ; Braunecker_2013 ; Vazifeh_2013 ; Yazdani_2014 ; Ruby_2015 ; Pawlak_2016 ; Kuster_2022 , Andreev spin qubits Andreev_sq_th ; Park_2017 ; Mishra_2021 ; Andreev_sq_exp ; Vidal_2023 ; Spectr_Spin_Andreec_SC ; Hoffman_2024 , Caroli-de Gennes-Matricorn and Majorana vortex states Volovik_1999 ; SSAR2_ex ; Caroli1964 ; Xu2015 ; Chiu2020 etc. Furthermore, our results offer insights for designing devices functioning as Cooper pair splitters and for realizing poor man’s Majorana fermions in minimal Kitaev chain models in which precise tunability over ECT and CAR processes is crucial Tunable_car_PRX ; MKC_ex ; MKC_ex_3dots . In our simulations, we employ realistic parameters. It is noteworthy that the coupling of side-leads to nanowires has already been experimentally demonstrated, both for normal leads topological_gap , for quantum dots Lee_Nat_Nan_2014 ; spin_filter_delftNC ; spin_filter_dots_Marcus , and, notably, for ferromagnetic leads Andreas_spin_ent_N . Hence, our predictions are directly amenable to experimental verification.
@@ II Model
In order to investigate the transport properties of the system we calculate numerically matrix elements of the differential-conductance $G_{ij}=dI_{i}/dV_{j}$ defined as derivative of the total current $I_{i}$ flowing into the nanowire from the lead $i$ with respect to the voltage bias $V_{j}$ applied to the lead $j$ as described in detail in Refs. Nonlocal_BCS_charge_th ; Hess_2021 , while the superconductor is grounded. To be specific we employ Blonder-Tinkham-Klapwijk formalism BTK in which the zero-temperature nonlocal conductance $G_{ij}(E)$ between leads $i$ and $j$ is given by
@@
The corresponding band structure of the leads $E_{L,R}(k)$ as a function of momentum $k$ is schematically depicted in Fig. 2 together with the band structure of the proximitized Rashba nanowire $E_{RNW}(k)$ and quasiparticle spin $S_{x}(k)=\Phi^{\dagger}(k)\tau_{z}\sigma_{x}\Phi(k)$ (see App. A for details). Again, the spin quantization axis is assumed to be in the $x$ direction, along the applied magnetic field. The leads are tunnel-coupled to the nanowire at the positions $x_{L}$ and $x_{R}$, respectively, for the left and the right lead with positive hopping amplitude $t_{\Gamma}(<t)$.
Figure 2: Energy bands of the left spin up polarized lead $E_{L}(k)$, the proximitized Rashba nanowire $E_{RNW}(k)$ and the right lead which is either (a) spin-down or (b) spin-up polarized or (c) normal $E_{R}(k)$. The $x$ component (along B-field) of the spin polarization of quasiparticles from the Andreev band, $S_{x}(k)$, is indicated by the colorbar. Shaded gray areas mark an energy window $\Delta_{i}<E<\Delta_{ex}$ in which spin sensitive transport occurs. Among many possible scattering processes, we denote the dominant ones contributing to the nonlocal conductances (a) $G_{L(+)R(-)}$, (b) $G_{L(+)R(+)}$, and (c) $G_{L(+)R(N)}$, which are either CAR (negative conductance) or ECT (positive conductance) processes. The strength of the corresponding signal (high or low) is reflected in the line thickness. Empty circles denote holes while filled ones denote electrons and half-filled ones correspond to quasiparticles in the nanowire. We set the following parameters for the nanowire: $\mu=0$, $t=0.27$ meV, $\Delta_{sc}=0.25$ meV, $\Delta_{Z}=0.2$ meV (nontopological phase), $\alpha=50$ meVnm ($E_{\rm SO}=0.23$ meV). For the spin up (down) polarized leads, we set $\mu_{\rm lead}=0$ and $M_{x}=-0.3$ meV ($M_{x}=0.3$ meV) and, for the normal lead, we set $\mu_{\rm lead}=0.3$ meV and $M_{x}=0$.
@@ III.1 Schematic picture of nonlocal transport
We present first a schematic (physical) picture of the dominant processes contributing to the nonlocal conductance $G_{LR}$ for three different configurations of the leads, which is then followed by the presentation of exact numerical results in subsequent sections. Regarding the proximitized Rashba nanowire, the extended Andreev band quasiparticles with energies $E_{RNW}(k)<\Delta_{sc}$ have a certain sign of the spin if the exterior gap exceeds the interior one, $\Delta_{i}<\Delta_{ex}$, which for $\mu=0$ is satisfied if $0<\Delta_{Z}<2\Delta_{sc}$. In the topologically trivial state (here, $\Delta_{Z}<\Delta_{sc}$ for $\mu=0$), the quasiparticle states with the lowest positive energy around $k\approx 0$ have negative spin polarization $S_{x}(k)$ as illustrated in Fig. 2(a)-(c) (middle panels).
When the left lead is spin-up (+) and the right lead spin-down polarized (-), the spin of the injected charge carrier with positive energy matches the negative spin of quasiparticles in the proximitized region. However, due to spin mismatch it cannot enter the left lead as an electron and has to be transformed into hole which is spin-down polarized [see Fig. 2(a) for details]. This leads to a strong CAR dominated nonlocal conductance $G_{L(+)R(-)}$ signal characterized by a negative sign.
@@
Finally, we consider the case with the left lead being spin-up polarized while the right lead is normal (unpolarized). Due to spin degeneracy in the right lead, the electron can freely enter the proximitized region regardless of the quasiparticle spin polarization and enter the left lead as spin-down (up) hole (electron). This supports strong CAR (ECT) dominated nonlocal conductances for positive (negative) energies [see Fig. 2(c) for details]. In the following sections, we support the picture presented here by exact numerical simulations of the spin-dependent quantum transport.
@@ III.2 Quasiparticle spin detection
We start the discussion with presenting results from transport simulations for systems with two leads that have the same spin polarization either up (+) or down (-). We observe a strong positive nonlocal conductance $G_{L(\pm)R(\pm)}(\Delta_{Z},E_{F})=G_{R(\pm)L(\pm)}(\Delta_{Z},E_{F})$, if the spin of the probed ABS is the same as the spin of the electrons in the leads. In this case the transport is dominated by ECT processes. For example, when the leads are spin up
@@
Figure 6: The same as for Fig. 5 but now the chemical potential $\mu$ is varied and the Zeeman energy is kept fixed to $\Delta_{Z}=0.4$ meV. Here, we can see again that the transport is dominated by CAR processes (negative conductance), which can be controlled by tuning $\mu$ (e.g. via backgate).
(spin down) polarized one can observe strong positive nonlocal conductance signals as depicted in Fig. 3(b’) [Fig. 3(b”)] for energies corresponding to the lowest bulk states and values of the Zeeman energies for which corresponding quasiparticle states also have positive (negative) sign of the local spin density at the positions of the leads. The appearance of weak negative nonlocal conductance signals (CAR) is due to the presence of SOI in the nanowire that allows for normally forbidden spin-flip transport processes. On the other hand, by analyzing plots of the local conductance shown in Fig.  3(c’) and (c”), we notice that conductance peaks are more (less) pronounced when the spin polarization in the leads matches (does not match) the spin polarization of the probed quasiparticles. This observation is consistent with so-called SSAR processes SSAR ; SSAR_ ; SSAR2_ex ; SSAR2_th ; SSAR_Dom ; Jeon_science_2017 ; Jeon_prb_2018 ; Wang_prl_2021 . We obtain analogous results with ECT dominated signals when the chemical potential deviates from $\mu=0$ while $\Delta_{Z}$ is fixed (see Fig. 4). In this case, there are two TPTs at $\mu_{c}^{\pm}=\pm\sqrt{\Delta_{Z}^{2}-\Delta_{sc}^{2}}$. Still, we again can identify these TPTs by simply looking at the sign of the nonlocal conductances.
@@
We have also considered the case when the left lead is spin polarized and the right one is normal - spin unpolarized (spin degenerate). In such a setup, the nonlocal conductance $G_{L(\pm)R(N)}$ has a strong positive (negative) value when the polarization of the left lead is the same as (opposite to) the spin polarization of the low-energy quasiparticles from the Andreev band. Such gate configurations make the system an ideal platform for detecting the sign of the local spin polarization of quasiparticles, which is directly linked to the sign of the nonlocal conductance. Therefore, such setups are very suitable for the detection of spin inversion of the one-dimensional Andreev bands induced by the TPT. Furthermore, such setups are optimal for switching between ECT and CAR processes in Rashba nanowires either by tuning the magnetic field [see Fig. 7(a,b)] or the chemical potential [see Fig. 8(a,b)]. For this case, we have also calculated $G_{L(\pm)R(N)}(x_{L})$ as a function of the position of the left lead while the position of the right one is fixed. We plot it together with quasiparticle spin and charge densities for the system in the topological as well as nontopological phase, see Fig. 9. We note again that the sign of the nonlocal conductance is correlated with the sign of the local spin density, which holds across a wide range of lead attachment positions. We also note that when the leads are attached to the ends of the system, the sign of the nonlocal conductance does not change when the system undergoes a TPT since the local spin density has the same sign at the ends of the nanowire (see Fig. 9 for details) in both topological and nontopological phases. Furthermore, we calculate nonlocal conductance maps $G_{L(\pm)R(N)}(\Delta_{Z},E)$ (see Fig. 10) in case where the leads are attached to the opposite ends of the nanowire. For such a setup one can observe that the sign of the nonlocal conductance in the trivial ($\Delta_{Z}<\Delta_{sc}$) and in the topological ($\Delta_{Z}>\Delta_{sc}$) phase is the same. This highlights the importance of side coupling of the leads to the regions that are away from the nanowire ends. Here, on purpose, we choose shorter nanowires $L=3\ \mu$m to illustrate clearly the case when the sign of the quasiparticle spin density $S_{x}(x)$ is position dependent [see orange curve on Fig. 9(b)]. The observed spin density oscillations and spin density sign change around the end of the nanowire, which occurs only in the topological phase, is determined essentially by the spin-orbit strength $\alpha$, Zeeman energy $\Delta_{Z}$ (more precisely by the induced gap $2|\Delta_{sc}-\Delta_{Z}|$), and the nanowire length $L$ which all affect the amplitude and frequency of oscillations and corresponding lengthscales (see App. D for details). As a consequence, in this case the sign of the nonlocal conductances $G_{L(\pm)R(N)}(x_{L})$ depends on the position of the left lead [see solid and dashed black curves on Fig. 9(b)]. We expect that such a position dependent nonlocal conductance measurement of the quasiparticle spin (charge) density could be more easily achieved experimentally in 2D systems hosting spin polarized YSR states or vortex states using a pair of spin polarized (normal) STM tips.
Figure 10: Maps of nonlocal conductance $G_{L(+)R(N)}(\Delta_{Z},E)$ in the case when the leads are attached at the opposite ends of the nanowire. Results shown in panel (a) are obtained for the same parameters as in Fig. 7 with $L=10$ $\mu$m and $t_{\Gamma}=0.5t\approx 0.14$ meV, while panel (b) illustrates results obtained for parameters from Fig. 9 with $L=3$ $\mu$m $t_{\Gamma}=0.5t_{c}\approx 1.5$ meV. Importantly, one can see clearly that the sign of the nonlocal conductance is the same for $\Delta_{Z}<\Delta_{sc}$ and for $\Delta_{Z}>\Delta_{sc}$. Thus, it is independent of the spin density of the inner region of the system and consequently of the topological phase of the system.
@@
Finally, we show that our results are robust against disorder, by adding random on-site fluctuations to the onsite chemical potential $\mu_{i}=\mu+\delta\mu_{i}$ in Eq. (1) with $|\delta\mu_{i}|\leq\Delta_{sc}$ as shown in Fig. 11(d) with corresponding mean free path $l_{mfp}~{}\simeq~{}685$ nm (see App. C for details). One can see that the presence of disorder affects the energy levels [see Fig. 11(a)] compared to the clean system [see Fig. 3(a)], however, the sign of the local spin density is very robust to such disorder [see Fig 11 (a,e)]. Moreover, as a consequence, our transport study shows that the sign of the nonlocal conductance is also not affected by the disorder [see Fig. 5(a, b’), Fig. 6(a,b’), Fig. 5(a,b’), Fig. 11(b,b’,b”,c,c’,c”)]. Importantly, in the charge neutrality regime where $\mu=0$ and where the band inversion is driven by the Zeeman energy, the spin-dependent nonlocal conductances are clearly more robust to the onsite disorder [see Fig. 5(a’,b’), Fig. 11(b,b’, b”,c,c’,c”)] than in the case when $\mu\neq 0$ [see Fig. 6(a’,b’)].